# On Minimal Achievable Quotas in Multiwinner Voting

### v1

@@ 4.1 General Properties
This approach generates a whole class of proportionality axioms. We, therefore, start by showing some general relational results: The known implications that EJR+ implies EJR and EJR implies JR only hold for $\alpha_{1}$-EJR+, $\alpha_{2}$-EJR, and $\alpha_{3}$-JR when $\alpha_{1}\leq\alpha_{2}\leq\alpha_{3}$. A complete proof, as well as all omitted proofs, can be found in the appendix.
@@
Let $\alpha_{1}\leq\alpha_{2}$ and $\Phi\in\{\text{JR, EJR, EJR+}\}$. First of all, $\alpha_{1}$-$\Phi$ implies $\alpha_{2}$-$\Phi$ since all cohesive groups for $\alpha_{2}$ are also considered for $\alpha_{1}$. Showing the implications between $\alpha$-EJR+, $\alpha$-EJR, and $\alpha$-JR follows directly from their definitions. ∎
@@
For $\alpha_{1}<\alpha_{2}<\alpha_{3}\leq 1$, it holds that $\alpha_{3}$-EJR+, $\alpha_{2}$-EJR, and $\alpha_{1}$-JR are incomparable.
Remark 2.
For $\alpha=0$, every voter group qualifies as cohesive, so no committee can satisfy the proportionality axioms; hence, they are all equivalent in this degenerate case.
We next demonstrate that there can be a substantial gap between the optimal $\alpha$-values for JR and EJR. This highlights that our framework provides greater flexibility in how representation is understood. Under the classical definitions, it is natural to focus on EJR, as it always implies JR. However, introducing the parameter $\alpha$ allows us to adjust the strength of representation: we can choose whether to prioritize smaller, cohesive groups or to emphasize the representation of larger ones.
@@
The additive gap between JR and EJR for the optimal $\alpha$-value can be $k/(k+1)$.
@@
The committee $W=\{c_{1},\dots,c_{k-1},c_{k+1}\}$ achieves $\alpha_{\text{JR}}(W)=0$, while for $\alpha_{\text{EJR}}$ we need to decide whether to select candidates $c_{1},\dots,c_{k}$ or to replace on of them with $c_{k+1}$. Either we get a violation by a $1$-cohesive group or by a $k$-cohesive group. When we set $x=ky$ in Figure˜3, we can show that $\alpha_{\text{EJR}}(W^{\prime})=k/(k+1)$ for all $W^{\prime}$. ∎
This construction shows that even when a committee perfectly satisfies $\alpha$-JR ($\alpha=0$), its $\alpha$-EJR value can be as large as $k/(k+1)$, demonstrating a maximal additive gap. Since $\alpha$ can be any value in $\mathbb{R}_{\geq 0}$, we prove that it is sufficient to only consider $k\cdot n$ many $\alpha$-values to determine the optimum for JR and EJR.
@@
For a given instance $I$, at most $\lceil\frac{n}{k}\rceil-1\leq n$ $\alpha$-values need to be checked to determine $\alpha^{*}_{\text{JR}}$. Further, it holds that $\alpha^{*}_{\text{JR}}\in\operatorname{\mathbb{Q}}$.
@@
For $\alpha$-JR, it is only relevant what size the smallest $1$-cohesive group has. This group size is between $1$, meaning that a single voter on its own already forms a $1$-cohesive group (together with a candidate she approves) and $\frac{n}{k}$ the size of the Hare quota. The relevant $\alpha$-values come from the set $X=\{0,\frac{k}{n},\frac{2k}{n},\dots,\frac{rk}{n}\}$ with $r=\lceil\frac{n}{k}\rceil-1$. Observe that if there does not even exist a violation for $\alpha=\frac{k}{n}$ then $\alpha^{*}_{\text{JR}}$ by definition is 0.
@@
For a given instance $I$, at most $n\cdot k$ many $\alpha$-values to determine the $\alpha_{\text{EJR}}^{*}$. Further,it holds that $\alpha_{\text{EJR}}^{*}\in\operatorname{\mathbb{Q}}$.
The last two lemmas allow us to assume that $\alpha\in\operatorname{\mathbb{Q}}$ since for any $\alpha\in\operatorname{\mathbb{R}}\setminus\operatorname{\mathbb{Q}}$ we can consider the next larger $\alpha^{\prime}$ from Lemma˜4.10 or Lemma˜4.11 respectively because if there exist a violation for $\alpha$, then there also exist one for $\alpha^{\prime}$.
@@ 4.2 Worst-Case Analysis of Existing Rules
Many known rules can always guarantee the Droop quota, meaning that $\alpha_{\text{JR}}^{*}(I)\leq\frac{k}{k+1}$ for every instance $I$. Nevertheless, there exist instances where the optimal $\alpha$-value is smaller by a factor of $\frac{n}{k+1}$. Intuitively, this follows from the fact that many rules greedily select candidates that are approved by a lot of voters and do not consider whether the voters who are underrepresented approve common candidates. Figure˜4 visualizes such an instance for $k=5$.
@@
Consider the following family of instances. Let $N=[(k+1)(k+1)]$. Furthermore, let $C=B\cup D$, with $B=\{b_{1},\dots,b_{k+1}\}$ such that $N_{b_{i}}=\{(i-1)\cdot(k+1)+1,\dots,i\cdot(k+1)\}$ and $D=\{d_{1},\dots,d_{k}\}$ such that $N_{d_{i}}=\{j\in N\colon(j\mod k+1)=i\}$ for all $i\in\{1,\dots,k+1\}$. Then each candidate in $D$ and each candidate in $B$ is approved by $k+1$ voters (see Figure˜4 for a visualization of the case $k=5$).
@@
Many existing voting rules achieve an $\alpha$-value of $\frac{k}{k+1}$ for this example since they can return the committee, which is visualized in violet, and leave the large group of voters $N_{b_{6}}$ unrepresented. Note that in this instance, rules like GJCR or MES would actually return not a single candidate since the instance has to contain a JR violation in order for them to select any candidate. We therefore consider a natural way to adapt these rules to $\alpha$ by lowering the quota and giving more budget to the voters, respectively. They are formally defined in Section˜4.2. In the following, we show that all these rules have a comparably bad $\alpha$-value on the instances defined in Example 4.12.
@@
For CC, seq-CC, seq-Phragmén, PAV, $\alpha$-MES, and $\alpha$-GJCR, the distance to the optimal $\alpha$-value can be of size $\frac{k^{2}}{(k+1)^{2}}$.
@@
Consider Example˜4.12: All of the above rules might choose $k$ from the $k+1$ candidates of the set $B$, i.e., the violet candidates from Figure˜4.
Without loss of generality, let $N_{b_{k+1}}$ be the group of voters that remain unrepresented. For them not to have a legitimate claim for a candidate in the committee, the quota $\alpha\cdot\frac{n}{k}$ must be greater than $\lvert N_{b_{k+1}}\rvert$. Consequently,
@@
On the other hand, note that there exists a committee $W=D$, such that for all pairs of non-represented voters $v,v^{\prime}$ with $v\neq v^{\prime}$ it holds that $A_{v}\cap A_{v^{\prime}}\cap A_{v^{\prime\prime}}=\emptyset$. Thus, it holds that $\alpha^{*}_{\text{JR}}(I)\leq\frac{k}{n}.$ Note that for this instance, $D$ is an optimal solution since there does not exist any committee that is approved by every voter. It follows that the $\alpha$-value of these voting rules can be $\frac{k}{k+1}-\frac{k}{n}=\frac{k^{2}}{(k+1)^{2}}$. ∎
Remark 3.
Example˜4.12 can be adapted so that it does not rely on tie-breaking. The ratio for this is then $\frac{n}{2(k+1)}$. (See example Example˜B.1 in Section˜B.1.)
@@
Theorem˜4.13 complements the results presented by Casey and Elkind (2025). They showed that the proportionality axioms defined with the Droop quota are still achievable by current voting rules. We show that these algorithms cannot provide any axiomatic guarantee in the sense of satisfying the more demanding $\alpha$-version of proportional representation notions when $\alpha$ is less than or equal to $\frac{k}{k+1}$ — corresponding to the Droop quota.
@@ 5 Computing the Optimal $\alpha$-value
For a given instance $I=(A,k)$ and committee $W$, we can efficiently compute $\alpha_{\text{JR}}(W)$.
@@
$s_{\max}=\max_{c\notin W}\lvert\{v\in N\colon A_{i}\cap W=\emptyset\text{ and }c\in A_{v}\}\rvert$.
@@
and thus, forms a violation with $c$. This implies that $\alpha_{\text{JR}}\geq s_{\max}\cdot\frac{k}{n}$.
On the other hand, by definition of $\alpha_{\text{JR}}$ there exists some $\alpha_{\text{JR}}$-JR violation with witness ($N^{\prime},c)$. Thus, it holds that $\lvert N^{\prime}\rvert\geq\alpha_{\text{JR}}\cdot\frac{n}{k}$. Further, since the voters in $N^{\prime}$ approve a common candidate $c$ and do not approve any candidate in $W$, it holds that
@@
Thus, $s_{\max}\geq\alpha_{\text{JR}}\cdot\frac{n}{k}$ which implies that $\alpha_{\text{JR}}\leq s_{\max}\cdot\frac{k}{n}$. ∎
Proposition˜5.1 proves that, as for JR, we can determine if there exists an $\alpha$-JR violation for a given committee in polynomial time. Nevertheless, finding $\alpha_{\text{JR}}^{*}(I)$ for an instance $I$, is NP-hard for all $0<\alpha<1$. The hardness also holds for EJR and EJR+. Remember that for $\alpha=0$ the proportionality axioms are always violated by definition, and for $\alpha>\frac{k}{k+1}$ they can always be satisfied.
@@
For all constant $0<\alpha<1$, deciding whether there exists a committee that satisfies $\alpha$-$\Phi$ with $\Phi\in\{\text{JR, EJR, EJR+}\}$ is NP-hard.
The intuition behind the reduction is as follows: We can ensure that certain candidates are selected if the support they receive is sufficiently large, and that every voter in the support only approves this specific candidate. Further, since $\alpha<1$, we can choose the support of those candidates small enough that after picking those candidates, there remains a large number of voters unrepresented. Avoiding an $\alpha$-JR violation for this group is then as hard as solving a vertex cover problem. The formal proof can be found in Section˜B.2.
Since the problem of determining the optimal $\alpha$-value for an instance is computationally hard, we provide a method to determine the value by repeatedly solving the following ILP:
$\sum_{c\in C}x_{c}$ $=k,\quad$ $\forall c\in C,\hskip 56.9055pt(1)$
$y_{v}$ $\leq\sum_{c\in A_{v}}x_{c}\quad$ $\forall v\in N,\hskip 56.9055pt(2)$
$\sum_{v:\,c\in A_{v}}(1-y_{v})\;$ $\leq\;\lceil\alpha\cdot\frac{n}{k}\rceil-1\quad$ $\forall c\in C,\hskip 56.9055pt(3)$
$y_{v},x_{c}\in\{0,1\}.$
@@
Condition (1) ensures that the committee contains precisely $k$ candidates, and condition (2) guarantees that, if $y_{v}=1$, the voter $v$ approves at least one candidate in the committee. The last condition ensures that at most $\lceil\alpha\cdot n/k\rceil-1$ voters approving some common candidate $c$ are left uncovered, capturing the $\alpha$-JR requirement. The constraint is equivalent to saying that at least $\lvert N_{c}\rvert-\lceil\alpha\cdot n/k\rceil+1$ voters in $N_{c}$ are covered. Finally, since we are only interested in the feasibility of the constraints, the ILP does not need an objective function. A proof of the correctness of the ILP formulation is established in Section˜B.2.
@@
For a fixed $\alpha$, a given instance $I$ and committee $W$, determining if $W$ satisfies $\alpha$-EJR is coNP-hard.
For a given instance $I$, figuring out the optimal $\alpha$-value with respect to EJR, i.e. finding $\alpha_{\text{EJR}}^{*}(I)$, can be solved by considering $\ell$-cohesive groups individually with $1\leq\ell\leq k$:
Namely, we can find for every $\ell$ the largest group of voters $s_{\max}(W,\ell)$ such that
@@
Similar to JR, we can now directly determine the largest $\alpha_{\ell}$ such that there exists an $(\alpha,\ell)$-cohesive group that causes an $\alpha$-EJR violation. Now for every $\ell$, we compute $\alpha_{\ell}=s_{\max}(W,\ell)\cdot\frac{k}{\ell\cdot n}$ and take the maximum over these values. This gives the largest $\alpha$-value such that there exists an $\alpha$-EJR violation. The reason that this does not give an efficient algorithm lies in the fact that determining $s_{\max}(W,\ell)$ is computationally hard for larger $\ell$. This computational complexity does not carry over to $\alpha$-EJR+.
@@
For a given instance $I=(A,k)$ and committee $W$, we can efficiently compute $\alpha_{\text{EJR+}}(W)$.

### v2

@@ 4.1 General Properties
This approach generates a whole class of proportionality axioms. We, therefore, start by showing some general relational results: The known implications chain of EJR+, EJR, and JR only holds for $\alpha_{1}$-EJR+, $\alpha_{2}$-EJR, and $\alpha_{3}$-JR when $\alpha_{1}\leq\alpha_{2}\leq\alpha_{3}$. Note that all proofs can be found in the full version of the paper.
@@
Let $\alpha\leq\alpha^{\prime}$ and $\Phi\in\{\text{JR, EJR, EJR+}\}$. First, $\alpha$-$\Phi$ implies $\alpha^{\prime}$-$\Phi$ since as $\alpha$ increases, the set of cohesive groups shrinks. Showing the implications between $\alpha^{\prime}$-EJR+, $\alpha^{\prime}$-EJR, and $\alpha^{\prime}$-JR then follows directly from the original definitions of the proportionality axioms. ∎
@@
For $1\geq\alpha_{1}>\alpha_{2}>\alpha_{3}$, it holds that $\alpha_{1}$-EJR+, $\alpha_{2}$-EJR, and $\alpha_{3}$-JR are incomparable.
We next demonstrate that there can be a substantial gap between the optimal $\alpha$-values for JR and EJR. This highlights that our framework offers greater flexibility in how representation is understood. Under the classical definitions, it is natural to focus on EJR, as it always implies JR. However, introducing the parameter $\alpha$ allows us to adjust the strength of representation: we can choose whether to prioritize smaller, cohesive groups or to emphasize the representation of larger ones.
@@
There is an instance $I$ such that the additive difference of $\alpha^{*}_{\text{JR}}(I)$ and $\alpha^{*}_{\text{EJR}}(I)$ is $k/(k+1)$.
@@
The committee $W=\{c_{1},\dots,c_{k-1},c_{k+1}\}$ achieves $\alpha_{\text{JR}}^{W}(I)=0$, while for $\alpha^{*}_{\text{EJR}}$ we need to decide whether to select candidates $c_{1},\dots,c_{k}$ or to replace one of them with $c_{k+1}$. Either we get a violation by a $1$-cohesive group or by a $k$-cohesive group. When we set $x=ky$ in Figure˜3, we can show that $\alpha^{*}_{\text{EJR}}=k/(k+1)$. ∎
This construction shows that even when a committee perfectly satisfies $\alpha$-JR ($\alpha=0$), its $\alpha$-EJR value can be as large as $k/(k+1)$, demonstrating a maximal additive gap. Since $\alpha$ can be any value in $\mathbb{R}_{\geq 0}$, we prove that it is sufficient to only consider $k\cdot n$ many $\alpha$-values to determine the optimum for JR and EJR. This technicality is important when designing efficient algorithms and is utilized in the restricted domain section.
@@
For a given instance $I$, at most $\lceil\frac{n}{k}\rceil-1(\leq n)$ $\alpha$-values need to be checked to determine $\alpha^{*}_{\text{JR}}(I)$. Further, it holds that $\alpha^{*}_{\text{JR}}(I)\in\operatorname{\mathbb{Q}}$.
@@
For $\alpha$-JR, it is only relevant what size the smallest $1$-cohesive group has. The group size is between $1$, meaning that a single voter forms a $1$-cohesive group (together with a candidate she approves), and $\frac{n}{k}$, the size of the Hare quota. The relevant $\alpha$-values come from the set $X=\{0,\frac{k}{n},\frac{2k}{n},\dots,\frac{rk}{n}\}$ with $r=\lceil\frac{n}{k}\rceil-1$. Observe that if there does not exist a violation for $\alpha=\frac{k}{n}$ then $\alpha^{*}_{\text{JR}}$ is 0 by definition.
@@
For a given instance $I$, at most $n\cdot k$ many $\alpha$-values are needed to determine $\alpha_{\text{EJR}}^{*}(I)$. Further, it holds that $\alpha_{\text{EJR}}^{*}(I)\in\operatorname{\mathbb{Q}}$.
The last two lemmas allow us to assume that $\alpha\in\operatorname{\mathbb{Q}}$ since for any $\alpha\in\operatorname{\mathbb{R}}\setminus\operatorname{\mathbb{Q}}$ we know from the previous two lemmas that there is some larger $\alpha^{\prime}\in\operatorname{\mathbb{Q}}$.
@@ 4.2 Worst-Case Analysis of Existing Rules
Many known rules can always guarantee the Droop quota. Nevertheless, there are instances in which these voting rules perform very poorly with respect to the optimal $\alpha$-value. Intuitively, this follows from the fact that many rules greedily select candidates approved by many voters and do not consider whether underrepresented voters approve of common candidates. Figure˜4 visualizes such an instance for $k=5$.
@@
Consider the following family of instances. Let $N=(k+1)(k+1)$. Furthermore, let $C=B\cup D$, with $B=\{b_{1},\dots,b_{k+1}\}$ such that $N_{b_{i}}=\{(i-1)\cdot(k+1)+1,\dots,i\cdot(k+1)\}$ for all $i\in[k+1]$ and $D=\{d_{1},\dots,d_{k}\}$ such that $N_{d_{i}}=\{j\in N\colon(j\mod k+1)=i\}$ for all $i\in[k]$. Then each candidate in $D$ and each candidate in $B$ is approved by $k+1$ voters (see Figure˜4 for a visualization of the case $k=5$).
@@
Many existing voting rules achieve an $\alpha$-value of $\frac{k}{k+1}$ in this example, since they return the committee, visualized in violet, while leaving the large group of voters $N_{b_{6}}$ unrepresented. Note that in this instance, rules like GJCR or MES would actually return no candidate, since the instance must contain a JR-violation for them to select any candidate. Therefore, we consider a natural way to adapt these rules to $\alpha$ by lowering the quota and giving more budget to the voters, respectively. We discuss these adaptations and define all the voting rules mentioned in the extended version. In the following, we show that all these rules have a comparably bad $\alpha$-value on the instances defined in Example˜4.12.
@@
For CC, seq-CC, seq-Phragmén, PAV, $\alpha$-MES, and $\alpha$-GJCR, the additive difference to $\alpha^{*}_{\text{JR}}(I)$ can be of size $\frac{k^{2}}{(k+1)^{2}}$.
@@
Consider Example˜4.12  [^4] Example 4.12 can be adapted so that it does not rely on tie-breaking. The additive difference for this is then $\frac{k^{2}}{(k+2)(k+1)}$ (see full version of the paper).: All of the above rules might choose $k$ from the $k+1$ candidates of the set $B$, i.e., the violet candidates from Figure˜4. Without loss of generality, let $N_{b_{k+1}}$ be the group of voters that remain unrepresented. For them not to have a legitimate claim for a candidate in the committee, the quota $\alpha\cdot\frac{n}{k}$ must be greater than $\lvert N_{b_{k+1}}\rvert$. Consequently,
@@
On the other hand, note that there exists a committee $W=D$, such that for all pairs of non-represented voters $v,v^{\prime}$ with $v\neq v^{\prime}$ it holds that $A_{v}\cap A_{v^{\prime}}=\emptyset$. Thus, it holds that $\alpha^{*}_{\text{JR}}(I)\leq\frac{k}{n}.$ Note that, in this instance, $D$ is an optimal solution since no committee is approved by every voter. It follows that the $\alpha$-value of these voting rules can be $\frac{k}{k+1}-\frac{k}{n}=\frac{k^{2}}{(k+1)^{2}}$. ∎
@@
Theorem˜4.13 complements the results presented by Casey and Elkind (2025). They demonstrated that the proportionality axioms defined with the Droop quota are still attainable under current voting rules. We show that these algorithms cannot provide any axiomatic guarantee in the sense of satisfying the more demanding $\alpha$-version of proportional representation notions when $\alpha$ is less than or equal to $\frac{k}{k+1}$ — corresponding to the Droop quota.
@@ 5 Computing the Optimal $\alpha$-value
For a given instance $I$ and committee $W$, we can compute $\alpha_{\text{JR}}^{W}(I)$ in $O(nm)$.
@@
$s_{\max}=\max_{c\notin W}\lvert\{v\in N\colon A_{v}\cap W=\emptyset\text{ and }c\in A_{v}\}\rvert$.
@@
and thus constitutes a violation of $c$. This implies that $\alpha_{\text{JR}}^{W}\geq s_{\max}\cdot\frac{k}{n}$.
On the other hand, by definition of $\alpha_{\text{JR}}^{W}$ there exists some $\alpha_{\text{JR}}^{W}$-JR violation with witness ($N^{\prime},c)$. Thus, it holds that $\lvert N^{\prime}\rvert\geq\alpha_{\text{JR}}^{W}\cdot\frac{n}{k}$. Further, since the voters in $N^{\prime}$ approve a common candidate $c$ and do not approve any candidate in $W$, it holds that
@@
Thus, $s_{\max}\geq\alpha_{\text{JR}}^{W}\cdot\frac{n}{k}$ which implies that $\alpha_{\text{JR}}^{W}\leq s_{\max}\cdot\frac{k}{n}$.
Since we need to compute $s_{\max}$, we must check, for every candidate $c\notin W$, how many voters are not represented and approve $c$. This can be done in $O(nm)$. ∎
Proposition˜5.1 proves that, as for JR, we can determine if there exists an $\alpha$-JR violation for a given committee in polynomial time.
@@
For a constant $\alpha\in(0,1)\cap\operatorname{\mathbb{Q}}$, deciding whether there exists a committee that satisfies $\alpha$-$\Phi$ with $\Phi\in\{\text{JR, EJR, EJR+}\}$ is NP-hard.
NP-completeness of $\alpha$-JR and $\alpha$-EJR+ then follows with Proposition˜5.1 and Proposition˜5.5, respectively. The intuition behind the reduction is as follows: We can ensure that certain candidates are selected if the support they receive is sufficiently large and every voter in that support approves only this specific candidate. Further, since $\alpha<1$, we can choose the support of those candidates small enough that after picking those candidates, there remains a large number of voters unrepresented. Avoiding an $\alpha$-JR violation for this group is then as hard as solving a vertex cover problem.
Since the problem of determining the optimal $\alpha$-value for an instance is computationally hard, we provide a method to determine the value by repeatedly solving the following ILP: [^5] Note that Lemma 4.10 limits the number of $\alpha$-values that need to be considered for the ILP.
$\sum_{c\in C}x_{c}=k,\quad\quad\quad\quad\;\;$ $\hskip 28.45274pt(1)$
$y_{v}\leq\sum_{c\in A_{v}}x_{c}\quad\forall v\in N,$ $\hskip 28.45274pt(2)$
$\sum_{v:\,c\in A_{v}}(1-y_{v})\;\leq\;\lceil\alpha\cdot\frac{n}{k}\rceil-1\quad\forall c\in C,$ $\hskip 28.45274pt(3)$
$y_{v},x_{c}\in\{0,1\}.\hskip 28.45274pt$
@@
Condition (1) ensures that the committee contains precisely $k$ candidates, and condition (2) guarantees that, if $y_{v}=1$, the voter $v$ approves at least one candidate in the committee. The last condition ensures that at most $\lceil\alpha\cdot n/k\rceil-1$ voters approving some common candidate $c$ are left uncovered, capturing the $\alpha$-JR requirement. The constraint is equivalent to saying that at least $\lvert N_{c}\rvert-\lceil\alpha\cdot n/k\rceil+1$ voters in $N_{c}$ are covered. Note that the ILP does not need an objective function, as we are only interested in feasibility.
@@
For a constant $\alpha\in(0,1]\cap\operatorname{\mathbb{Q}}$, a given instance $I$ and committee $W$, determining if $W$ satisfies $\alpha$-EJR is coNP-hard.
For a given instance $I$ and committee $W$, figuring out the optimal $\alpha$-value with respect to EJR, i.e. finding $\alpha_{\text{EJR}}^{W}(I)$, can be solved by considering $\ell$-cohesive groups individually with $1\leq\ell\leq k$:
Namely, we can find for every $\ell$ and $W$ the size of the largest group of voters, denoted $s_{\max}(W,\ell)$, such that
@@
Similar to JR, we can now directly determine the largest $\alpha_{\ell}$ such that there exists an $(\alpha,\ell)$-cohesive group that causes an $\alpha$-EJR violation. Now for every $\ell$, we compute $\alpha_{\ell}=s_{\max}(W,\ell)\cdot\frac{k}{\ell\cdot n}$ and take the maximum over these values. This gives the largest $\alpha$-value such that there exists an $\alpha$-EJR violation. The reason that this does not give an efficient algorithm lies in the fact that determining $s_{\max}(W,\ell)$ is computationally hard for larger $\ell$. This computational complexity does not carry over to $\alpha_{\text{EJR+}}^{W}(I)$.
@@
For a given instance $I$ and committee $W$, we can compute $\alpha^{W}_{\text{EJR+}}(I)$ in $O(mnk)$.